### v2

# Parameterized sequential functions for temporal properties

Abstract
Developers of “reactive” [4] software, like operating systems, still need to complete significant code development to have much idea about what performance and correctness problems or overlooked issues will force changes in the design or compromises in design goals. Temporal logic seemed to provide a way for developers to think about and validate these kinds of systems at a higher level, but progress has been slow at best. This paper shows that the kinds of properties that temporal logics promised to address, are among those that can be expressed in ordinary algebra with less notation and without the overhead of formal logic or axiomatic methods, using sequential functions[14] to define ordinary state machines.
1 Introduction
Everybody who has worked in formal logic will confirm that it is one of the technically most refractory parts of mathematics. — Von Neumann[8]
Temporal logic[11, 10, 7, 3, 5, 6] is based on an intuitively appealing claim that being able to specify when a property must be true can be useful in specifications of computer systems. Examples include properties like: “Always no more than one process can be accessing a mutually exclusive resource”, and “Eventually every scheduled process must start to run”, and “a server cannot send an acknowledgment Until it has received the request.” This paper shows how to express similar and more detailed properties without formal logics or other axiomatic system, using sequential functions [14]. A sequential function is a map from finite sequences of events to outputs so that $f(w)$ is the output of a discrete state system in the state reached by following event sequence $w$ from the initial system state. As seen below, these functions can specify detailed operation of large scale state systems as well as abstract properties and the behavior and architecture of composite and concurrent systems.
Section 2 starts with a short introduction to sequential functions and a definition of temporal-logic-style operators on boolean valued sequential functions so that $(\mathop{\mathit{Always}}P)(w)$, $(\mathop{\mathit{Eventually}}P)(w)$ and so on are well-defined and can be seen to be at least similar to the corresponding temporal qualifiers (which differ among themselves in any case). Section 3 tackles a number of examples, including real-time and concurrent systems, and contrasts specifications that do and do not use the temporal-logic-like operators. Section 4 covers a real-time example used in a paper by Abadi and Lamport[1] and includes a detailed proof. The final section tries for some context.
The goal of this project is perhaps anachronistic: it is to provide human developers of operating systems and other complex “reactive” software [4] with a mathematical basis for specifying and thinking about designs. The reader can form their own opinion about how successful this effort has been and whether it is more or less successful than temporal logic.
2 Basic methods and Temporal-like operators
Each sequential function is associated with an alphabet – the set of discrete events that can can produce a state change. Every state change is due to some event in the alphabet, although it is common practice to not completely list the alphabet as shown below. State systems are deterministic, so any finite sequence of events determines the state reached by following the sequence from the initial state. All event sequences in this paper are finite.
If $w$ is a event sequence and $f$ is a sequential function, then $f(w)$ is a output in the state reached by following $w$ from the initial state. Following the zero length sequence $\epsilon$ from the initial state leaves the system state unchanged so $f(\epsilon)$ is the output in the initial state. Appending an event $a$ to a sequence, $w\cdot a$, drives the system to the next state so $f(w\cdot a)$ is the output in the state reached after event $a$ drives the system from the state determined by $w$. By convention $\epsilon$ is not an element of any alphabet and $w\cdot\epsilon=w$ leaving the system in the same state (see section 3.3 for why this is convenient).
If $z$ is a finite sequence, then $w\mathop{\mathtt{concat}}z$ concatenates $z$ to $w$ on the right and $f(w\mathop{\mathtt{concat}}z)$ is the output in the state reached by following $z$ from the state determined by $w$. Primitive recursion on sequences completely defines a sequential function:
$f(\epsilon)=c\text{ for constant }c,\ f(w\cdot a)=g(a,f(w)).$
Simple composition
$f_{2}(w)=h(f_{1}(w))$
modifies the output and compositions of multiple sequential functions e.g.
$f(w)=(f_{1}(w),\dots f_{n}(w))$
define direct products of state systems (which do not interact). More complex products can be defined using sequential functions as transducers: see section 3.3.
Boolean functions have values in the set $\{0,1\}$ with $1$ for true and 0 for false. Boolean expressions like $f(w)<c$ will be treated as abbreviations for characteristic boolean functions like
$P(w)=\begin{cases}1&\text{if }f(w)<c\\ 0&\text{otherwise}\end{cases}$
Sequential functions depend on a single event sequence argument, which is usually given as the first argument. Multiple arguments can mean that the output set $X$ consists of maps (generally not sequential functions), so $f(w,x,y,z)=f(w)(x,y,z)$ or it may be that $f(w)=(f_{1}(w),f_{2}(w),\cdots f_{n}(w))$ and $f(w,i)=(f(w))_{i}=f_{i}(w)$.
More details on sequential functions can be found in appendix B and in the earlier paper [14] where it is shown that sequential functions are essentially just Moore-type state machines and products of these machines. The focus here is on intuition and applications not the underlying automata theory.
2.1 Temporal style operators
Concatenation of finite sequences, sequence prefixes, and sequence length are all well known operations, but definitions can be found in appendix A in case of doubt.
Definition 1
$(\mathop{\mathit{Always}}P)(w)=\forall z\text{ where }P(w\mathop{\mathtt{concat}}z)\text{ is defined, }P(w\mathop{\mathtt{concat}}z)$
[…]
$(P\mathop{\mathit{Until}}Q)(w)=\forall z\text{ where }P(w\mathop{\mathtt{concat}}z)\text{ is defined and }P(w\mathop{\mathtt{concat}}z)=0$
$Q(w\mathop{\mathtt{concat}}z^{\prime})\text{ for some prefix }z^{\prime}\text{ of }z$
[…]
Note that if $P(w)$ then $(\mathop{\mathit{Eventually}}P)(w)$ with $n=0$. ‘ The treatment of undefined states above is not necessarily the best for every application. It might also to be good to require that neither next nor eventually are trivially true, with at least one $a$ so that $P(w\cdot a)$ is defined and one sufficiently long sequence $z$ so that $P(w\mathop{\mathtt{concat}}z)$ is defined.
2.1.1 Some Conventions
An expression $(\mathop{\mathit{Eventually}}P(w,x))$ has the meaning $(\mathop{\mathit{Eventually}}P_{x})(w)$ where $P_{x}(w)=P(w)(x)$
Nested temporal style operators sometimes require a little care – like nested summations or integrals do. Consider $(\mathop{\mathit{Always}}(\mathop{\mathit{Eventually}}P))(w)$. Let $Q(w)=(\mathop{\mathit{Eventually}}P)(w)$. Then the original expression can be written $(\mathop{\mathit{Always}}Q)(w)$ which means that for all $z$ so that $Q(w\mathop{\mathtt{concat}}z)$ is defined, $Q(w\mathop{\mathtt{concat}}z)$. This last expression is defined to mean
$(\mathop{\mathit{Eventually}}P)(w\mathop{\mathtt{concat}}z)$
which means that there is some $n$ so that if $u$ has length $n$ or more, for some prefix $u^{\prime}$ of $u$,
$P(w\mathop{\mathtt{concat}}z\mathop{\mathtt{concat}}u^{\prime}).$
Suppose $Sent(w,i,m)$ is true iff network element $i$ has sent message $m$, $Received(w,i,m)$ has the obvious meaning and $Sending(w,i,m)$ is true if $i$ is currently sending the message. For these, suppose that if $Sent(w,i,m)>Sent(w\cdot a,i,m)$ is impossible so that a message never becomes unsent, and the same for $Received$. Then
$\text{If }Sending(w,i,m)\text{ then }(\mathop{\mathit{Eventually}}Sent)(w,i,m)$ (1)
would be a convenient property. If message $m$ must be acknowledged with message $m^{\prime}$ and
$\begin{array}[]{l}\text{If }Sent(w,i,m)<Received(w,i,m)\\ \text{ then }(\mathop{\mathit{Eventually}}Sending)(w,i,m))\end{array}$ (2)
requires that element $i$ keep sending the message until it gets the acknowledgment.
Alternatively, define (SinceP)(w)(Since\ P)(w) to be the count of events since $P$ first became true.
$(Since\ P)(\epsilon)=0$
$(Since\ P)(w\cdot a)=\begin{cases}P(w)&\text{if }(Since\ P)(w)=0\\ 1+(Since\ P)(w)&\text{otherwise}\end{cases}$
Then for some $n\geq 0$ (SinceSending)(w,i,m))(1Sent(w,i,m)))<n(Since\ Sending)(w,i,m))*(1-Sent(w,i,m)))<n is equivalent to 1 above and for some $n_{1}\geq 0$ $((SinceSent)(w,i,m))*(1-Received(w,i,m^{\prime}))<n_{1}$ for some fixed $n_{1}$, is the same constraint as in 2 above. These constraints assume that there are no errors, which is not always a good assumption. It may be better to allow the network element to conclude that a client has failed or is unreachable if for some $k$ (SinceSent)(w,i,m)>k(Since\ Sent)(w,i,m)>k and $Received(w,i,m)=0$. Section 3.4 shows how to count elapsed time instead of the number of events.
3 Examples
3.0.1 Simple examples
Some simple examples illustrate and will be useful in section 4. The common theme is to define state recursively. The functions “tailtail” and “headhead used below are defined in appendix A. used below.
Specification 1
Components.
  • $S$ is a store over values $V$ if $S(w\cdot v)=v$ for every $v\in V$. It is a store over values $V$ with initial value $v_{0}$ if $S(\epsilon)=v_{0}$. Otherwise the initial value is not specified.
  • $T$ is a toggle with initial value $b\in\{0,1\}$ if $T(\epsilon)=b$ and $T(w\cdot a)=1-T(w)$.
  • $Q$ is a transparent queue over set $V$ if $Q(\epsilon)=\epsilon$ and
    $Q(w\cdot a)=\begin{cases}Q(w)\cdot v&\text{if }a=(enq,v)\\ tail(Q(w))&\text{if }a=deq\text{ and }Q\neq\epsilon\\ Q(w)&\text{otherwise}\end{cases}$
  • $Q$ is a bounded queue over set $V$ with length $k$ if $Q(\epsilon)=\epsilon$ and:
    $Q(w\cdot a)=\begin{cases}Q\cdot v&\text{if }a=(enq,v)\text{ and }Length(Q(w))<k\\ tail(Q)v&\text{if }a=deq\text{ and }Q\neq\epsilon\\ Q(w)&\text{otherwise}\end{cases}$
  • $Q$ is a closed queue over set $V$ with length $k$ if for some bounded queue with length $k$, $Q_{k}$ over $V$
    $Q(w)=\begin{cases}\epsilon&\text{if }Q_{k}(w)=\epsilon;;\\ head(Q_{k}(w))&\text{otherwise}\end{cases}$
    Here, $\epsilon\not\in V$ is assumed.
3.1 Process scheduling
Figure 1: State graph of process scheduling status
Suppose $Status(w)\in\{free,ready,running,blocked\}$ is the scheduling status of a process within an operating system and the event alphabet is $A_{sched}=\{schedule,wake,preempt,run,block,free\}$. Scheduling status can be represented by a simple state graph as in figure 1 or with the following sequential function.
[…]
The only difference between this state machine and the one defined by the state graph is the unrecoverable error state on an unexpected event. We could just forbid those events with $Status(w)\neq error$, but that constraint would then be on the wrong level because the schedule component does not control its inputs.
To track which core a process is run on replace the single runrun event with $(run,c)$ for $c\in Cores$ where $Cores$ is a set of processor core identifiers. To track the reason a process is blocked, do the same thing with $(block,r)$ and a set of $Reasons$.
$A_{sched}=\{schedule,wake,preempt,(run,c)\ c\in Cores,(block,r)\ r\in Reasons,free\}.$
Then change the function definition.
[…]
Even this change would make the graph in figure 1 impractical to draw if either the set of core names or the set of reasons had more than a very few elements.
The specification does not satisfy:
$\text{If }Status(w)=ready\text{ then }(\mathop{\mathit{Eventually}}Status(w)=(running,c))$
for some $c\in Cores$. The schedule component depends on the operating system for liveness as seen in section 3.3. The techniques of that section would also permit interconnection of other sequential functions that capture other process behavior. For example, $Req(w)\in Requests\cup\{null\}$ might capture the current request to the operating system, if any.
3.2 Clock pulse
Figure 2: Specifying a clock signal
Suppose the system events are binary $n$-tuples that represent discrete time samples of input signals at a fixed frequency so $(a)_{i}$ is the signal on wire $i$ carried by event (sample) $a$. Because the frequency is fixed, counting events tracks the passage of time. Say $C$ is a circuit with $m$ output signals if $C(w)\in\{0,1\}^{m}$. Let $(C(w))_{i}$ be the $i^{th}$ element of that output tuple. Suppose output $k$ is supposed to be a clock signal with a cycle constrained by time parameters $t_{0},t_{1}$ where $t_{1}>t_{0}$ and the signal is supposed to stay stable until at least $t_{0}$ time units and then switch levels before $t_{1}$ time units have passed. See figure 2. What happens to the other output signals or what effects the input signals have on them is not specified, yet. Define $Stable(w,k,b)$ to count the time duration that output $k$ has been stable at value $b\in\{0,1\}$.
$\begin{array}[]{l}Stable(\epsilon,k,b)=0\\ Stable(w\cdot a,k,b)=((C(w\cdot a))_{k}=b)*(1+Stable(w,k,b))\end{array}$
Then the desired circuit property could be written:
$\text{If }C(w)_{k}=b\text{ then }C(w)_{k}=b\mathop{\mathit{Until}}Stable(w,k,b)\geq t_{0}$
and
$Stable(w,k,b)<t_{1}$
or we could replace the statement using $\mathop{\mathit{Until}}$ with
$\text{If }C(w)_{k}\neq C(w\cdot a)_{k}\text{ then }Stable(w,k,b)\geq t_{0}$
3.3 Multiple processes
Suppose the schedule state systems from section 3.1 are components of an operating system or network system with its own event alphabet and there is set $ProcessIds$ of process identifiers. For $p\in ProcessIds$ a set of process identifiers, let:
$Pstatus(w,p)=Status(u_{p}(w))$
where $u_{p}$ is a connector sequential function that translates sequences of the enclosing system into sequences over the component alphabet $A_{sched}$. When the system is in the state determined by $w$, schedule component $p$ is in the state determined by $u_{p}(w)$. The constraint $Pstatus(w,p)\neq error$ requires the connector functions to not generate events that the schedulers don’t expect.
Connector function definitions have two standard forms. The simpler case, where components advance by one or zero events on every event of the enclosing system looks like this:
$u_{p}(\epsilon)=\epsilon,\text{ and }u_{p}(w\cdot a)=u_{p}(w)\cdot g_{p}(w,a).$
The equation $u_{p}(\epsilon)=\epsilon$ puts the components in their initial states when the operating system is in its initial state. The initial value can be any constant sequence over the component alphabet. The map $g_{p}$ is called the tail function and $g_{p}(w,a)\in A_{sched}\cup\{\epsilon\}$ When the system advances by event $a$ from the state determined by $w$, schedule component $p$ advances by $g_{p}(w,a)$ or remains in its current state if $(g_{p}(w,a)=\epsilon)$. The more general form components can also advance by multiple steps on a single event for the enclosing state system (see appendix B).
There can be many different kinds of components in an interconnected system and figure 3 is a partial illustration of the operating system with components. For the $Req$ sequential function of section 3.1 there might be connector maps $v_{p}$ for each $p\in Processes$ and
$\begin{array}[]{l}g_{p}(w,a)=(block,r)\text{ only if }Req(v_{p}(w))=r\text{ and}\\ \text{ for some }c,\ Sched(u_{p}(w))=(running,c)\end{array}$
Figure 3: Multiple process state machines, interconnected
A requirement that no core is asked to run more than one process in any state constrains $u_{p}$.
[…]
This property could be made more concrete with:
$\begin{array}[]{l}\text{If }g_{p}(w,a)=(run,c)\text{ then }\Sigma_{p}(Pstatus(w,p)=(running,c))=0\\ \text{ and }\Sigma_{p}(g_{p}(w,a)=(run,c))=1\\ \text{ and }Pstatus(w,p)=ready\end{array}$
Now it makes sense to require scheduling be live.
$\begin{array}[]{l}\text{If }Pstatus(w,p)=ready\text{ then }\\ (\mathop{\mathit{Eventually}}Pstatus(w,p)=(running,c))\text{ for some }c\end{array}$
This property could be a useful first approximation for example to show some network algorithm eventually gets to the right place. On the other hand, general purpose time-shared operating system generally cannot guarantee this property. We’d need at least an admissions policy to limit the number of active processes and additional controls on memory and other resources. A crude requirement to keep the number of active processes below the number of cores.
$\Sigma_{p}(Pstatus(w,p)\neq free)\leq|Cores|$
would be interesting only for very simple real-time systems. Attempting to specify how a general purpose operating system could guarantee liveness or at least identify the failure conditions that would break the guarantee, is a good problem. More realistically we could define a boolean sequence function to test whether timing specifications have been met, Then qualifying any propositions about the system with with this test limits those propositions to states where the timing specification has been met.
3.4 Real-time process constraints
The enclosing system can contain additional components along with processes. One component could be a real-time clock, which may not correspond to an actual device but could just be an ideal clock that somehow extracts time from events at the enclosing system level.
$Clock(w\cdot a)>Clock(w)$ (3)
In a deterministic event driven system, the passage of real-time must be represented by events since nothing else changes states but not all components need to change state as time advances. Which events cause time to pass can differ by what physical system is being represented. For the circuit above described by figure 2, events are real-time samples of signal values so every event marks passage of one unit of time. In the operating system or network context, the passage of time might be marked by clock ticks from an oscillator, or clock interrupts, or some kind of input value like a time received from a GPS satellite or a vector of samples of signals on all input wires of the top level system during a discrete moment of time. It is also possible to take a real valued time variable $t$ and just require that
$Clock(w)\sim t\ .$ (4)
That choice doesn’t have to be made at this point. Let
$Elapsed(w,a)=Clock(w\cdot a)-Clock(w).$ (5)
Now it is possible to track how long a process has been runnable, waiting to run. Define $Delay(w,p)$ as follows:
$Delay(\epsilon,p)=0$
$Delay(w\cdot a,p)=\begin{cases}Delay(w,p)+Elapsed(w,a)&\text{if }Pstatus(w\cdot a,p)=ready\\ &\text{and }Pstatus(w,p)=ready\\ 0&\text{if }Pstatus(w\cdot a,p)\neq ready\\ Delay(w,p)&\text{otherwise}\end{cases}$
ClockClock depends on $w$ the sequence driving the operating system or network directly, while $Pstatus(w,p)$ depends on $u_{p}(w)$.
Define $RTSched(w,x)$ for $x>0$ by $RTSched(\epsilon,x)=1$ and
$RTSched(w\cdot a)=\begin{cases}0&\text{if for some }p,\ Delay(w\cdot a,p)>x\\ RTSched(w,x)&\text{otherwise}\end{cases}$
To require that every scheduled process must eventually run:
$\text{for some }x_{0}>0,RTSched(w,x_{0})$
It is certainly possible to imagine the bound being state dependent, so
$Delay(w,p)\leq Bound(w).$
BoundBound might depend on how many processes are waiting to run and what resources they need and their priorities.
The constraint using the $\mathop{\mathit{Eventually}}$ operator is not the same as any of these, because the eventually operator counts each event as one time unit and keeps the count invisible.
4 A Message queue
Figure 4: The queue from Abadi/Lamport
A paper by Abadi and Lamport [1] includes an example of a somewhat peculiar buffered queue. The queue is first specified in a way that makes it unreliable (it can lose input data) and then it is fixed by adding real-time constraints. I’ve tried to stay close to the approach in the Abadi and Lamport paper which they summarize as follows:
The interface consists of two pairs of “wires”, each pair consisting of a val wire that holds a message and a boolean-valued bit wire. A message m is sent over a pair of wires by setting the val wire to m and complementing the bit wire. The receiver detects the presence of a new message by observing that the bit wire has changed value. Input to the queue arrives on the wire pair (ival, ibit), and output is sent on the wire pair (oval, obit). There is no acknowledgment protocol, so inputs are lost if they arrive faster than the queue processes them. (Because of the way ibit is used, inputs are lost in pairs.) The property guaranteed by this lossy queue is that the sequence of output messages is a subsequence of the sequence of input messages [^1] This is not correct. VY. [Below] we add timing constraints to rule out the possibility of lost messages. (page 1545)
The core of the pre-realtime Abadi and Lamport specification is in figure 5 but the reader here doesn’t need to examine it to understand this specification.
The motivating example might be a A/D device or simple input device that has a single datum (message) input buffer, a program or device that is responsible for collecting and queuing the messages, and a third component that dequeues messages and makes them available to a consumer in the output buffer. The queue is lossy because is is possible that messages that are stored to the input buffer can be overwritten by a second inputinput event before they are moved into the internal queue.
From the outside, ignoring the components for the moment, suppose there is an alphabet $A_{queue}$ which contains events including deqdeq and $(input,m)$ for every message $m\in Msgs$ and other events which can be left for later. The output is a pair of the message at the head of the queue and a bit that alternates on each message.
[…]
The output message is nullmnullm if the queue was empty on the last deqdeq.
Here, the buffered queue is constructed from the store, toggle, and queue components from specification 1 above. There are three toggles and two buffers and one transparent queue. A $(input,m)$ event stores message $m$ in the input buffer and toggles the input toggle. A deqdeq event causes the internal queue to move the head of the queue to the output buffer and toggles the output toggle. An internal enqenq event copies the message in the input buffer to the internal queue and flips the $Last$ toggle bit. The components of the queue and connector functions are given below, with component names starting with a capital letter (e.g. $Last)$ and connector functions all lower case: lastlast.
Specification 2
There is a set of messages $\mathit{Msgs}$. InvalInval and $Outval$ are stores over $\mathit{Msgs}$ with initial value nullmnullm; $Q$ is a transparent queue over $\mathit{Msgs}$; $Inb,Last$, and $Outb$ are toggles with initial value 0.
The connector maps are $ov,ob,inb,last,\text{ and }q: invinv$ for $Inval$, inbinb for $Inb$, ovov for $Outval$, obob for $Outb$, lastlast for $Last$, $q$ for $Q$.
The initial values of the connector functions are all $\epsilon$.
$inv(\epsilon)=ov(\epsilon)=inb(\epsilon)=last(\epsilon)=ob(\epsilon)=q(\epsilon)=\epsilon$
The toggles all start out with output 0 per the specification. The message buffers are both set to $\mathit{nullm}$ initially. For example, since $inv(\epsilon)=\epsilon$ and $Invar$ is a store with initial value nullmnullm, $Invar(inv(\epsilon))=nullm$.
Each component advances only one or zero steps for every composite system event in $A_{queue}$ since we let $w\cdot\epsilon=w$.
All connector maps have the form
$u(\epsilon)=c,\ u(w\cdot a)=u(w)\cdot{u}^{\mathsf{T}}(w,a)$
so specifying the recursive tail is most of what is needed. The convention that the tail of $u$ is named ${u}^{\mathsf{T}}$ is just a naming convention, the T doesn’t have any other meaning.
An $(input,m)$ event produces an event stores $m$ to the $Inval$ component and toggles the $Inb$ toggle. Use the convention that $(E\neq E^{\prime})$ has value one if the two expressions are not equal and 0 otherwise and then set
$WaitEnq(w)=(Line(line(w))\neq Inb(inb(w)))$
The intent of the protocol with the toggles is that $WaitEnq(w)=1$ when $Inval(iv(w))$ is an input message that has not yet been queued on the internal queue. The most complex connection map is $q$ which controls events for the input queue and which is not completely specified but is constrained from enqueuing unless the $Last$ and $Inb$ values differ. Incomplete specification takes the place of nondeterminism.
Specification 3
The recursive “tails” of the connectors.
[…]
Finally:
$Queue(w)=(Oval(ov(w)),Obit(ob(w)))$
4.1 Real-time
To patch up the poor design of the protocol in this queue, Abadi and Lamport propose to put timing constraints on input and enq operations. The basic idea here is that when an inputinput event happens, a timer starts and the matching enqenq must happen before $t_{0}$ seconds pass, but the next input event cannot happen until more than $t_{0}$ seconds have passed. It’s pretty obvious that these timing constraints fix the problem with dropped messages, but see section 4.2.
Let’s use the ClockClock described in equation 3 and 4 and ElapsedElapsed from equation 5 above.
[…]
SinceISinceI tracks the time that has passed since the last inputinput event. In the initial state, it is set to $t_{0}$ so that the initial inputinput can happen at any time without flagging an error.
$\begin{array}[]{l}SinceI(\epsilon)=t_{0}\\ SinceI(w\cdot a)=\begin{cases}0&\text{if }a=(input,m)\\ &\text{ for some }m\in\mathit{Msgs}\\ SinceI(w)+Elapsed(w,a)&\text{otherwise}\end{cases}\end{array}$
There cannot be an inputinput event without triggering a timing error until $SinceI(w)>t_{0}$. There must be an enqenq event before $SinceI(w)\geq t_{0}$.
Putting these together the timing requirement is:
[…]
Once $Tspec(w)=0$ it stays that way, because a single timing error is not recoverable for this queue.
The timing specification is not requiring that events follow the specification which could be done with $\mathop{\mathit{Always}}Tspec(w)=1$. Instead the specification allows us to only consider states that are reached without violations of the specification.
For comparison, see page 1553 of [1] which is similar at least in intent.
4.2 Correctness
We want to prove that if $Tspec(w)=1$ then the queue has not dropped or reordered or invented any queue elements. Abadi and Lamport characterize this property as something like, the sequence of input values is a prefix of the sequence of output values. We’re going to prove something stronger but first, let’s define what those two sequences are.
Define: $Q_{I}(\epsilon)=\epsilon$ and
$Q_{I}(w\cdot a)=\begin{cases}Q_{I}(w)\cdot m&\text{if }a=(input,m)\text{ for some }m\in\mathit{Msgs}\\ Q_{I}(w)&\text{otherwise}\end{cases}$
This is the abstract queue of inputs. Define: $Q_{O}(\epsilon)=\epsilon$ and
$Q_{O}(w\cdot a)=\begin{cases}Q_{O}(w)\cdot m&\text{if }{q}^{\mathsf{T}}(w,a)=(deq,m)\text{ for some }m\in\mathit{Msgs}\\ Q_{O}(w)&\text{otherwise}\end{cases}$
This is the queue of messages that has passed through the device to the output buffer. The queues are similar to the history variables of Abadi and Lamport. The theorem (it’s really too simple to be called a theorem) to prove is stronger than what Abadi and Lamport suggest.
Theorem 1
Assuming $Tspec(w)$
$\text{If }WaitEnq(w)=0\text{ then }Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))$
$\text{If }WaitEnq(w)=1\text{ then }Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))\cdot Inval(in(w))$
The difference is that when $WaitEnq(w)=1$ the value in the input buffer is the last item that has shown up on the input queue, otherwise the input buffer can be ignored. It’s an immediate consequence of this theorem that $Q_{O}(w)$ is a prefix of $Q_{I}(w)$.
The proof of 1 is by induction on the event sequence. Initially, when $w=\epsilon$, all the queues are empty and $WaitEnq(\epsilon)=0$ so there is nothing to prove. Suppose the theorem is true for $w$ and consider $w\cdot a$. We assume $Tspec(w\cdot a)$ otherwise there is nothing to prove. There are four cases.
  1. 1.
    case:
    $WaitEnq(w)=0\text{ and }WaitEnq(w\cdot a)=1$
    By the definition of $WaitEnq$, one of the two toggles has to change thanks to $a$. If ${q}^{\mathsf{T}}(w,a)=(deq,m)$ that would change the $Last$ toggle but because $WaitEnq(w)=0$ ${q}^{\mathsf{T}}(w,a)\neq(deq,m)$ for any $m$ so it must be that $a=(input,m)$ for some $m$ and $Inb$ changes. By the inductive assumption and because $WaitEnq(w)=0$ $Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))$. Because $a=(input,m)$, $Q_{I}(w\cdot a)=Q_{I}(w)\cdot m$. By the definition of ${inv}^{\mathsf{T}}$ and because $a=(input,m)$, $Inval(inv(w\cdot a))=m$. We know $a\neq deq$ so ${q}^{\mathsf{T}}(w,a)\neq deq$ – so $Q(q(w))=Q(q(w\cdot a))$. This proves the first case.
  2. 2.
    case:
    $WaitEnq(w)=0\text{ and }WaitEnq(w\cdot a)=0$
    In this case $a\neq(input,m)$ because otherwise $WaitEnq(w\cdot a)=1$ unless the $Last$ toggle flips too. But $Last$ cannot change because that needs ${q}^{\mathsf{T}}(w,a)=(enq,m)$ and that cannot be true if $WaitEnq(w)=0$. So $Q_{I}(w\cdot a)=Q_{I}(w))$. If $a=deq$ then if $Q(q(w))=\epsilon$ nothing happens to $Q_{O}$ or $Q$. If $Q(q(w))\neq\epsilon$ and $a=deq$, then ${q}^{\mathsf{T}}(w,a)=deq$ so $Q(q(w\cdot a)=Tail(Q(w))$ and $Q_{O}(w\cdot a)=Q_{O}(w)\cdot Head(Q(q(w)))$ but
    $Q_{O}(w)\cdot Head(Q(q(w))\mathop{\mathtt{concat}}Tail(Q(q(w))$
    must be the same as
    $Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w)).$
    Otherwise ${q}^{\mathsf{T}}(w,a)=\epsilon$ That takes care of this case.
  3. 3.
    case:
    $WaitEnq(w)=1\text{ and }WaitEnq(w\cdot a)=1$
    Because $WaitEnq$ remains true, ${q}^{\mathsf{T}}(w,a)\neq(enq,m)$ for any $m$ so no new elements are added to $Q(q(w\cdot a))$. Because we assume $Tspec(w\cdot a)$, $WaitEnq(w)*SinceI(w)<t_{0}$ so $a$ is not an input event so ${inv}^{\mathsf{T}}(w,a)=\epsilon$ and $Inval(inv(w\cdot a))=Inval(inv(w))$. If $a\neq deq$, it must be that ${q}^{\mathsf{T}}(w,a)=\epsilon$ so $Q(q(w\cdot a))=Q(q(w))$ when $a\neq deq$ and similarly $Q_{O}(w\cdot a)=Q_{O}(w)$ because ${ov}^{\mathsf{T}}(w,a)=\epsilon$. If $a=deq$ then than $Q(q(w\cdot a))=Tail(Q(q(w)))$ and $Q_{O}(w\cdot a)=Q_{O}(w)\cdot Head(Q(q(w)))$ so the constructed queue is unchanged.
    For any other $a$ nothing changes in the outputs of any components and we are done with this case.
  4. 4.
    case:
    $WaitEnq(w)=1\text{ and }WaitEnq(w\cdot a)=0$
    Since $WaitEnq(w)=1$
    $Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))\cdot Inv(inv(w).$
    Either $SinceI(w)<t_{0}$ or $TSpec(w)=0$. We’re assuming the second is false, so by the same reasoning $a\neq(input,m)$ (or else $Tspec(w\cdot a)=0$). That means that $q(w,a)=(enq,Inv(inv(w)))$ to make $WaitEnq(w\cdot a)=0$. As a result
    $Q(q(w\cdot a))=Q(q(w))\cdot Inv(inv(w)).$
    And the rest follows.
Figure 5: Abadi and Lamport’s queue specification (untimed)
4.3 Some notes
The original specification depends on an unbounded queue, but in this case imposing a bound $k$ on the queue could be accomplished by replacing the transparent queue with a bounded queue of length $k$ which would make an enq operation when the queue is full just discard the message in the in buffer but would still toggle the value in $Last$. Given a timing specification for deqdeq we could show that for some $k$ the queue would be long enough. An interesting way to specify timing would be to require that deq operations happen every $t_{1}$ seconds at the latest, but that once a deq operation happened, other deq’s would happen every $t_{2}$ for some $t_{2}<t_{1}$ until the queue had been emptied.
5 Discussion
When Professor Krithivasan Ramamritham gave me temporal logic papers to read[12] at the start of my graduate studies in the 1980s, I had just finished working on a commerical distributed operating system[2] and was frustrated by how often the developers had found design errors only during test (after writing enormous quantities of code). Temporal logic seemed like a solution, but I was not unique in finding that it was not a completely satisfactory answer. Treatment of composition and concurrency, the inflexibility of the temporal quantifiers, the ad hoc nature of the formal semantics were among the issues that made it difficult to get a handle on the complexity of actual operating system design. Much of the methods used here and in [14] were present in a 1991 paper [13], which included Moore machines, automata products for composition and concurrency and even a modal primitive recursive specification language. But that work was in the tradition of computational formal logic and this is not. When computer scientists started to try to develop mathematical methods of specifying programs, they often adopted methods from formal logic and other axiomatic methods perhaps because of the role of formal languages in programming languages and pattern matching and because of the influence of metamathematics on computability studies. Whatever the merits of that approach, discarding the apparatus of formal logic has allowed solution of a number of problems that blocked progress in the earlier method and greatly simplified specifications.
The quote from Von Neumann that starts this paper is from a, not surprisingly, enormously perceptive paper that among other things previews algorithmic complexity theory. The full paragraph in which that quote is found has an interesting critique which I have tried to learn from as I returned to this work after a long absence.
We are very far from possessing a theory of automata which deserves that name, that is, a properly mathematical-logical theory. There exists today a very elaborate system of formal logic, and, specifically, of logic as applied to mathematics. This is a discipline with many good sides, but also with certain serious weaknesses. This is not the occasion to enlarge upon the good sides, which I have certainly no intention to belittle. About the inadequacies, however, this may be said: Everybody who has worked in formal logic will confirm that it is one of the technically most refractory parts of mathematics. The reason for this is that it deals with rigid, all-or-none concepts, and has very little contact with the continuous concept of the real or of the complex number, that is, with mathematical analysis. Yet analysis is the technically most successful and best-elaborated part of mathematics. Thus formal logic is, by the nature of its approach, cut off from the best cultivated portions of mathematics, and forced onto the most difficult part of the mathematical terrain, into combinatorics.
Appendix A Appendix: Sequence utilities
$w\mathop{\mathtt{concat}}\epsilon=w,\ w\mathop{\mathtt{concat}}(z\cdot a)=(w\mathop{\mathtt{concat}}z)\cdot a)$
$length(\epsilon)=0,\ length(w\cdot a)=1+length(w)$
$prefixes(\epsilon)=\{\epsilon\},\ prefixes(w\cdot a)=prefixes(w)\cup\{w\cdot a\}$
$rappend(\epsilon,a)=a\cdot\epsilon,rappend(w\cdot b,a)=rappend(w,a)\cdot b)$
$Head(\epsilon)=\epsilon,Head(rappend(w,a))=a$
$Tail(\epsilon)=\epsilon,Tail(rappend(w,a))=w$
Appendix B Appendix: Sequential functions
This is a short summary of results from [14]. The standard presentation of a Moore machine is $(A,X,S,s_{0},\delta,\lambda)$ where $s_{0}\in S$ is the start state, $\delta:S\times A\to S$ and $\lambda:S\to X$. Primitive recursive sequential functions are reformulations to make the state set implicit and the recursion explicit.
$D(\epsilon)=c,\ D(w\cdot a)=g(D(w),a)\text{ and }M(w)=L(D(w))$
Such a map can be transformed into a conventional (possibly infinite state) Moore machine by letting the state set be the set $\{D(w):w\in A^{*}\}$.
A sequential function is any map $f:A^{*}\to X$ where $A$ is a set of events, $A^{*}$ are the finite sequences over $A$ and $X$ is the set of outputs. Sequential functions define possibly infinite state Moore machines via the usual congruence on sequences.
The primitive recursive sequence functions are a proper subset with some useful properties including $f(w)$ must be defined if $f(w\cdot a)$ is defined. The finite state sequential functions are a proper subset of the primitive recursive sequential functions that correspond to finite state Moore machines. Specifications of real devices and programs tend to use primitive recursive and finite state primitive recursive sequential functions without any special attention because properties of real systems are constructive and real systems are finite state machines anyway.
Primitive recursive sequential functions are constructed by a finite number of applications of the following rules.
  1. 1.
    Constant. If $c$ is a constant $f(w)=c$ defines a sequential function.
  2. 2.
    Constant map. If $g$ is a non-state dependent map, $f(w,x)=g(x)$ defines a sequential function.
  3. 3.
    Primitive Recursion on Sequences. If $g$ is a non-state dependent map, $c$ is a constant, $\epsilon$ is the empty sequence
    $f(\epsilon)=c,\ f(w\cdot a)=g(a,f(w)))$
    defines a sequential function. If $g$ has a finite image $f$ is finite state.
  4. 4.
    Simple Composition. If $f_{1}$ is a sequential function, and $h(x)$ is an ordinary map
    $f(w)=h(f_{1}(w))$
    defines a sequential function.
As with the arithmetic primitive recursive functions[9], many additional forms can be defined without changing the class of primitive recursive sequential functions. In particular multiple arguments $f(w,x_{1},\dots x_{n})$ are fine, even with some arguments that are also state dependent $f(w)=(f(v_{1}(w),\dots f_{n}(w)))$. An analog of arithmetic simultaneous recursion allows definition of composite system which correspond to automata products and specify concurrent/parallel systems of machines.
  • General Product
    Given sequential functions $f_{1},\dots f_{n}$, each $f_{i}:A_{i}^{*}\to X_{i}$ and appropriate maps $g_{i}:A\times X\to A_{i}^{*}$ (not sequence dependent) where $X=X_{1}\times\dots X_{n}$,
    $f(w)=(f_{1}(u_{1}(w)),\dots f_{n}(u_{n}(w)))$
    $u_{i}(\epsilon)=\epsilon,\text{ and }u_{i}(w\cdot a)=u_{i}(w)\mathop{\mathtt{concat}}g(a,i,v(w))$
As shown in [14]: If each of the $f_{i}$ is primitive recursive, then $f$ is primitive recursive. If each $f_{i}$ is finite state and $A$ is finite then $f$ is finite state.

### v3

@@
 Abstract
-Developers of “reactive” [4] software, like operating systems, still need to complete significant code development to have much idea about what performance and correctness problems or overlooked issues will force changes in the design or compromises in design goals.
-Temporal logic seemed to provide a way for developers to think about and validate these kinds of systems at a higher level, but progress has been slow at best.
-This paper shows that the kinds of properties that temporal logics promised to address, are among those that can be expressed in ordinary algebra with less notation and without the overhead of formal logic or axiomatic methods, using sequential functions[14] to define ordinary state machines.
+Temporal logic is based on an intuitively appealing claim that being able to specify when a property will be true can be useful in specifications of computer software and hardware - particularly in reactive systems [6].
+This paper shows that temporal properties are among the properties that can be expressed directly and more simply in terms of state machines using sequential functions[17] to define state machine behavior and compositional architecture.
+Specifications and proofs can be carried out in ordinary algebra without the need for a formal language or axiomatic proof system.
+The compositional properties of sequential functions also simplify specification of composite and concurrent systems.
 1 Introduction
@@
 Temporal logic[13, 12, 10, 3, 7, 9] is based on an intuitively appealing claim that being able to specify when a property must be true can be useful in specifications of computer systems.
-Examples include properties like: “Always no more than one process can be accessing a mutually exclusive resource”, and “Eventually every scheduled process must start to run”, and “a server cannot send an acknowledgment Until it has received the request.” This paper shows how to express similar and more detailed properties without formal logics or other axiomatic system, using sequential functions [14].
-A sequential function is a map from finite sequences of events to outputs so that $f(w)$ is the output of a discrete state system in the state reached by following event sequence $w$ from the initial system state.
+Examples include properties like: “Always no more than one process can be accessing a mutually exclusive resource”, and “Eventually every scheduled process must start to run”, and “a server cannot send an acknowledgment Until it has received the request.” This paper shows how to express similar and more detailed properties without formal logics or other axiomatic system using sequential functions [17].
+A sequential function is a map from finite sequences of events to outputs so that $f(w)$ is the output of a discrete state system in the state reached by following event sequence $w$ from the initial state.
 As seen below, these functions can specify detailed operation of large scale state systems as well as abstract properties and the behavior and architecture of composite and concurrent systems.
@@
 Section 3 tackles a number of examples, including real-time and concurrent systems, and contrasts specifications that do and do not use the temporal-logic-like operators.
-Section 4 covers a real-time example used in a paper by Abadi and Lamport[1] and includes a detailed proof.
+Section 4 covers a real-time example used in a paper by Abadi and Lamport[1] but unlike that paper includes a complete proof.
 The final section tries for some context.
 The goal of this project is perhaps anachronistic: it is to provide human developers of operating systems and other complex “reactive” software [6] with a mathematical basis for specifying and thinking about designs.
-The reader can form their own opinion about how successful this effort has been and whether it is more or less successful than temporal logic.
+The reader can form their own opinion about how successful this effort has been.
 2 Basic methods and Temporal-like operators
-Each sequential function is associated with an alphabet – the set of discrete events that can can produce a state change.
-Every state change is due to some event in the alphabet, although it is common practice to not completely list the alphabet as shown below.
+Each sequential function is associated with an alphabet – the set of discrete symbols representing the events that can can produce a state change.
 State systems are deterministic, so any finite sequence of events determines the state reached by following the sequence from the initial state.
-All event sequences in this paper are finite.
-If $w$ is a event sequence and $f$ is a sequential function, then $f(w)$ is a output in the state reached by following $w$ from the initial state.
-Following the zero length sequence $\epsilon$ from the initial state leaves the system state unchanged so $f(\epsilon)$ is the output in the initial state.
-Appending an event $a$ to a sequence, $w\cdot a$, drives the system to the next state so $f(w\cdot a)$ is the output in the state reached after event $a$ drives the system from the state determined by $w$.
-By convention $\epsilon$ is not an element of any alphabet and $w\cdot\epsilon=w$ leaving the system in the same state (see section 3.3 for why this is convenient).
+Event sequences in this paper are always finite.
+Much of the expressive power of sequential functions derives from how convenient it is to make partial specifications: instead of making nondeterminism an intrinsic property, we can just leave some behavior unspecified.
+Following the zero length sequence $\epsilon$ from the initial state leaves the system state unchanged so for sequential function $f$, $f(\epsilon)$ is the output in the initial state.
+Appending an event $a$ to a sequence, $w\cdot a$, drives the system to the next state so $f(w\cdot a)$ is the output in the state reached after event $a$ drives the system one event from the state determined by $w$.
 If $z$ is a finite sequence, then $w\mathop{\mathtt{concat}}z$ concatenates $z$ to $w$ on the right and $f(w\mathop{\mathtt{concat}}z)$ is the output in the state reached by following $z$ from the state determined by $w$.
 Primitive recursion on sequences completely defines a sequential function:
-$f(\epsilon)=c\text{ for constant }c,\ f(w\cdot a)=g(a,f(w)).$
+ $\text{ for constant }c,\ f(\epsilon)=c,\text{ and }f(w\cdot a)=g(a,f(w)).$
 Simple composition
 $f_{2}(w)=h(f_{1}(w))$
-modifies the output and compositions of multiple sequential functions e.g.
+modifies the output. Compositions of multiple sequential functions e.g.
 $f(w)=(f_{1}(w),\dots f_{n}(w))$
-define direct products of state systems (which do not interact). More complex products can be defined using sequential functions as transducers: see section 3.3.
-Boolean functions have values in the set $\{0,1\}$ with $1$ for true and 0 for false. Boolean expressions like $f(w)<c$ will be treated as abbreviations for characteristic boolean functions like
+define direct products of state systems (which do not interact). An analog of simultaneous recursion of arithmetic primitive recursive functions [11] defines a system of interacting components:
+$f(w)=(f_{1}(w_{1}),\dots f_{n}(w_{n}))$
+where each $w_{i}=w_{i}(w)$ is a primitive recursive sequential function that determines component interactions. See section 3.3 for details.
+Boolean valued sequential functions that have values in the set $\{0,1\}$ with $1$ for true and 0 for false can be used to specify properties.
+Boolean expressions like $f(w)<c$ will be treated as abbreviations for characteristic boolean functions like
 $P(w)=\begin{cases}1&\text{if }f(w)<c\\ 0&\text{otherwise}\end{cases}$
 Sequential functions depend on a single event sequence argument, which is usually given as the first argument.
-Multiple arguments can mean that the output set $X$ consists of maps (generally not sequential functions), so $f(w,x,y,z)=f(w)(x,y,z)$ or it may be that $f(w)=(f_{1}(w),f_{2}(w),\cdots f_{n}(w))$ and $f(w,i)=(f(w))_{i}=f_{i}(w)$.
-More details on sequential functions can be found in appendix B and in the earlier paper [14] where it is shown that sequential functions are essentially just Moore-type state machines and products of these machines.
+Multiple arguments can mean either that that the output set $X$ consists of maps (generally not sequential functions), so $f(w,x,y,z)=(f(w))(x,y,z)$.
 The focus here is on intuition and applications not the underlying automata theory.
 2.1 Temporal style operators
-Concatenation of finite sequences, sequence prefixes, and sequence length are all well known operations, but definitions can be found in appendix A in case of doubt.
 Definition 1
@@
 […]
+Sequence prefixes and sequence length are well known operations, but definitions can be found in appendix A in case of doubt.
 Note that if $P(w)$ then $(\mathop{\mathit{Eventually}}P)(w)$ with $n=0$.
@@
 Then the original expression can be written $(\mathop{\mathit{Always}}Q)(w)$ which means that for all $z$ so that $Q(w\mathop{\mathtt{concat}}z)$ is defined, $Q(w\mathop{\mathtt{concat}}z)$.
-This last expression is defined to mean
+Expanding $Q$ the last expression means
 $(\mathop{\mathit{Eventually}}P)(w\mathop{\mathtt{concat}}z)$
@@
 Suppose $Sent(w,i,m)$ is true iff network element $i$ has sent message $m$, $Received(w,i,m)$ has the obvious meaning and $Sending(w,i,m)$ is true if $i$ is currently sending the message.
-For these, suppose that if $Sent(w,i,m)>Sent(w\cdot a,i,m)$ is impossible so that a message never becomes unsent, and the same for $Received$.
+For these, suppose that $Sent(w,i,m)>Sent(w\cdot a,i,m)$ is impossible so that a message never becomes unsent, and the same for $Received$.
 Then
@@
 would be a convenient property. If message $m$ must be acknowledged with message $m^{\prime}$ and
-$\begin{array}[]{l}\text{If }Sent(w,i,m)<Received(w,i,m)\\ \text{ then }(\mathop{\mathit{Eventually}}Sending)(w,i,m))\end{array}$ (2)
+$\begin{array}[]{l}\text{If }Sent(w,i,m)>Received(w,i,m^{\prime})\\ \text{ then }(\mathop{\mathit{Eventually}}Sending)(w,i,m))\end{array}$ (2)
 requires that element $i$ keep sending the message until it gets the acknowledgment.
@@
 $(Since\ P)(w\cdot a)=\begin{cases}P(w)&\text{if }(Since\ P)(w)=0\\ 1+(Since\ P)(w)&\text{otherwise}\end{cases}$
-Then for some $n\geq 0$ (SinceSending)(w,i,m))(1Sent(w,i,m)))<n(Since\ Sending)(w,i,m))*(1-Sent(w,i,m)))<n is equivalent to 1 above and for some $n_{1}\geq 0$ $((SinceSent)(w,i,m))*(1-Received(w,i,m^{\prime}))<n_{1}$ for some fixed $n_{1}$, is the same constraint as in 2 above.
+Then for some $n\geq 0$, (SinceSending)(w,i,m))(1Sent(w,i,m)))<n(Since\ Sending)(w,i,m))*(1-Sent(w,i,m)))<n is equivalent to 1 above and for some $n_{1}\geq 0$ $((SinceSent)(w,i,m))*(1-Received(w,i,m^{\prime}))<n_{1}$ for some fixed $n_{1}$, is the same constraint as in 2 above.
 These constraints assume that there are no errors, which is not always a good assumption.
@@
 The common theme is to define state recursively.
-The functions “tailtail” and “headhead used below are defined in appendix A.
-used below.
+The functions “tailtail” and “headhead used below are as usual and are defined in appendix A in case of doubt.
 Specification 1
@@
 Suppose the schedule state systems from section 3.1 are components of an operating system or network system with its own event alphabet and there is set $ProcessIds$ of process identifiers.
-For $p\in ProcessIds$ a set of process identifiers, let:
+For $p\in ProcessIds$, let:
 $Pstatus(w,p)=Status(u_{p}(w))$
@@
 The constraint $Pstatus(w,p)\neq error$ requires the connector functions to not generate events that the schedulers don’t expect.
-Connector function definitions have two standard forms. The simpler case, where components advance by one or zero events on every event of the enclosing system looks like this:
-$u_{p}(\epsilon)=\epsilon,\text{ and }u_{p}(w\cdot a)=u_{p}(w)\cdot g_{p}(w,a).$
-The equation $u_{p}(\epsilon)=\epsilon$ puts the components in their initial states when the operating system is in its initial state.
+Connector functions are defined recursively:
+$u_{p}(\epsilon)=\epsilon,\text{ and }u_{p}(w\cdot a)=u_{p}(w)\mathop{\mathtt{concat}}G_{p}(w,a).$
+The equation $u_{p}(\epsilon)=\epsilon$ puts the process components in their initial states when the operating system is in its initial state.
 The initial value can be any constant sequence over the component alphabet.
-The map $g_{p}$ is called the tail function and $g_{p}(w,a)\in A_{sched}\cup\{\epsilon\}$ When the system advances by event $a$ from the state determined by $w$, schedule component $p$ advances by $g_{p}(w,a)$ or remains in its current state if $(g_{p}(w,a)=\epsilon)$.
-The more general form components can also advance by multiple steps on a single event for the enclosing state system (see appendix B).
+The map $G_{p}$ is called the tail function and $G_{p}(w,a)$ is what happens to the component $p$ when the system advances by an $a$ event from the system state determined by $w$.
+If $G_{p}(w,a)=\epsilon$, component $p$ remains in the same state.
+For this simple example suppose each process component only advances 0 or $1$ steps on each step of the enclosing system and write $a$ for $\langle a\rangle$ if it is clear from the context that the single element sequence being used.
 There can be many different kinds of components in an interconnected system and figure 3 is a partial illustration of the operating system with components.
-For the $Req$ sequential function of section 3.1 there might be connector maps $v_{p}$ for each $p\in Processes$ and
+For the $Req$ sequential function of section 3.1 there might be connector maps $V_{p}$ for each $p\in Processes$.
 $\begin{array}[]{l}G_{p}(w,a)={(block,r)}\text{ only if }Req(V_{p}(w))=r\text{ and}\\ \text{ for some }c,\ Sched(u_{p}(w))=(running,c)\end{array}$
@@
 […]
-This property could be made more concrete with:
-$\begin{array}[]{l}\text{If }g_{p}(w,a)=(run,c)\text{ then }\Sigma_{p}(Pstatus(w,p)=(running,c))=0\\ \text{ and }\Sigma_{p}(g_{p}(w,a)=(run,c))=1\\ \text{ and }Pstatus(w,p)=ready\end{array}$
+This property could be made more direct with:
+$\begin{array}[]{l}\text{If }G_{p}(w,a)={(run,c)}\text{ then }Pstatus(w,p)=ready\\ \text{ and }\Sigma_{p}(Pstatus(w,p)=(running,c))=0\\ \text{ and }\Sigma_{p}(G_{p}(w,a)={(run,c))}=1\end{array}$
 Now it makes sense to require scheduling be live.
@@
 $Delay(\epsilon,p)=0$
-$Delay(w\cdot a,p)=\begin{cases}Delay(w,p)+Elapsed(w,a)&\text{if }Pstatus(w\cdot a,p)=ready\\ &\text{and }Pstatus(w,p)=ready\\ 0&\text{if }Pstatus(w\cdot a,p)\neq ready\\ Delay(w,p)&\text{otherwise}\end{cases}$
+$Delay(w\cdot a,p)=\begin{cases}Delay(w,p)+Elapsed(w,a)&\text{if }Pstatus(w\cdot a,p)=ready\\ &\text{and }Pstatus(w,p)=ready\\ 0&\text{otherwise}\end{cases}$
 ClockClock depends on $w$ the sequence driving the operating system or network directly, while $Pstatus(w,p)$ depends on $u_{p}(w)$.
@@
 To require that every scheduled process must eventually run:
-$\text{for some }x_{0}>0,RTSched(w,x_{0})$
+$\text{for some }x,\ RTSched(w,x)$
 It is certainly possible to imagine the bound being state dependent, so
@@
 The queue is lossy because is is possible that messages that are stored to the input buffer can be overwritten by a second inputinput event before they are moved into the internal queue.
-From the outside, ignoring the components for the moment, suppose there is an alphabet $A_{queue}$ which contains events including deqdeq and $(input,m)$ for every message $m\in Msgs$ and other events which can be left for later.
+From the outside, ignoring the components for the moment, suppose there is an alphabet $A_{queue}$ which contains events including deqdeq and $(input,m)$ for every message $m\in\mathit{Msgs}$ and other events which can be left unspecified.
 The output is a pair of the message at the head of the queue and a bit that alternates on each message.
@@
 The output message is nullmnullm if the queue was empty on the last deqdeq.
-Here, the buffered queue is constructed from the store, toggle, and queue components from specification 1 above.
-There are three toggles and two buffers and one transparent queue.
+The queue is constructed from the store, toggle, and queue components from specification 1 above.
+Specification 2
+$\mathit{Msgs}$ is a set of messages.
+InvalInval and $Outval$ are stores over $\mathit{Msgs}$ with initial value nullmnullm;
+InbInb, OutbOutb, and $Last$ are toggles with the same initial value.
+$Q$ is a transparent queue over $\mathit{Msgs}$;
 A $(input,m)$ event stores message $m$ in the input buffer and toggles the input toggle.
@@
 An internal enqenq event copies the message in the input buffer to the internal queue and flips the $Last$ toggle bit.
-The components of the queue and connector functions are given below, with component names starting with a capital letter (e.g.
-$Last)$ and connector functions all lower case: lastlast.
-Specification 2
-There is a set of messages $\mathit{Msgs}$.
-InvalInval and $Outval$ are stores over $\mathit{Msgs}$ with initial value nullmnullm; $Q$ is a transparent queue over $\mathit{Msgs}$; $Inb,Last$, and $Outb$ are toggles with initial value 0.
-The connector maps are $ov,ob,inb,last,\text{ and }q: invinv$ for $Inval$, inbinb for $Inb$, ovov for $Outval$, obob for $Outb$, lastlast for $Last$, $q$ for $Q$.
-The initial values of the connector functions are all $\epsilon$.
-$inv(\epsilon)=ov(\epsilon)=inb(\epsilon)=last(\epsilon)=ob(\epsilon)=q(\epsilon)=\epsilon$
-The toggles all start out with output 0 per the specification.
-The message buffers are both set to $\mathit{nullm}$ initially.
-For example, since $inv(\epsilon)=\epsilon$ and $Invar$ is a store with initial value nullmnullm, $Invar(inv(\epsilon))=nullm$.
-Each component advances only one or zero steps for every composite system event in $A_{queue}$ since we let $w\cdot\epsilon=w$.
-All connector maps have the form
-$u(\epsilon)=c,\ u(w\cdot a)=u(w)\cdot{u}^{\mathsf{T}}(w,a)$
-so specifying the recursive tail is most of what is needed. The convention that the tail of $u$ is named ${u}^{\mathsf{T}}$ is just a naming convention, the T doesn’t have any other meaning.
-An $(input,m)$ event produces an event stores $m$ to the $Inval$ component and toggles the $Inb$ toggle.
-Use the convention that $(E\neq E^{\prime})$ has value one if the two expressions are not equal and 0 otherwise and then set
-$WaitEnq(w)=(Line(line(w))\neq Inb(inb(w)))$
-The intent of the protocol with the toggles is that $WaitEnq(w)=1$ when $Inval(iv(w))$ is an input message that has not yet been queued on the internal queue.
-The most complex connection map is $q$ which controls events for the input queue and which is not completely specified but is constrained from enqueuing unless the $Last$ and $Inb$ values differ.
-Incomplete specification takes the place of nondeterminism.
+The intent of the protocol with the toggles is that the $Inb$ and $Last$ toggles are not equal if and only if there is a live message in $Inval$ tht has not yet been enqued in the internal queue.
+The connector maps are $w_{1}$ for $Inb$ and $Inval$, $w_{2}$ for the toggle $Last$, $w_{3}$ for the internal queue $Q$, and $w_{4}$ for the output bit toggle and output buffer.
+Say $w_{i}(w\cdot a)$ adds an event $a$ if and only if $w_{i}(w\cdot a)=w_{i}(w)\cdot a$.
+Say $w_{i}(w\cdot a)$ adds nothing if and only if $w_{i}(w\cdot a)=w_{i}(w)$.
+The specification for $w_{3}$ is different from the others because we don’t want to fully specify when $w_{3}(w\cdot a)$ adds an enqenq event.
+In the Abadi and Lamport specification this is a nondeterministic choice, here it depends on the unspecified variations among the components, the other input events, and possibly other factors not specified.
+Section 4.1 adds real-time constraints.
 Specification 3
-The recursive “tails” of the connectors.
 […]
-Finally:
-$Queue(w)=(Oval(ov(w)),Obit(ob(w)))$
+The queue device itself then just selects from the tuple value of $RawDevice$.
+Specification 4
+$Queue(w)=((RawDevice(w))_{5},(RawDevice(w))_{6})$
 4.1 Real-time
 To patch up the poor design of the protocol in this queue, Abadi and Lamport propose to put timing constraints on input and enq operations.
-The basic idea here is that when an inputinput event happens, a timer starts and the matching enqenq must happen before $t_{0}$ seconds pass, but the next input event cannot happen until more than $t_{0}$ seconds have passed.
+The basic idea here is that when an $(input,m)$ event happens, a timer starts and the matching enqenq must happen before $t_{0}$ seconds pass, but the next input event cannot happen until more than $t_{0}$ seconds have passed.
 It’s pretty obvious that these timing constraints fix the problem with dropped messages, but see section 4.2.
@@
 $\begin{array}[]{l}SinceI(\epsilon)=t_{0}\\ SinceI(w\cdot a)=\begin{cases}0&\text{if }a=(input,m)\\ &\text{ for some }m\in\mathit{Msgs}\\ SinceI(w)+Elapsed(w,a)&\text{otherwise}\end{cases}\end{array}$
-There cannot be an inputinput event without triggering a timing error until $SinceI(w)>t_{0}$. There must be an enqenq event before $SinceI(w)\geq t_{0}$.
+There cannot be an $(input,m)$ event without triggering a timing error until $SinceI(w)>t_{0}$. But also once there is an input event, there must be an enqenq event before $SinceI(w)\geq t_{0}$.
+To make the specification more compact: $WaitEnq$ indicates if there is a live message in the inbuffer according to the protocol.
+$WaitEnq(w)=\begin{cases}1&\mbox{ if }Last(w_{2}(w)))\neq Inb(w_{1}(w)))\\ 0&\text{otherwise}\end{cases}$
 Putting these together the timing requirement is:
@@
 The timing specification is not requiring that events follow the specification which could be done with $\mathop{\mathit{Always}}Tspec(w)=1$.
-Instead the specification allows us to only consider states that are reached without violations of the specification.
-For comparison, see page 1553 of [1] which is similar at least in intent.
+Instead the specification allows us to distinguish states that are reached without violations of the specification.
+Those reached otherwise have unspecified results.
+For comparison, see page 1553 of [1] which is similar in intent.
 4.2 Correctness
@@
 This is the abstract queue of inputs. Define: $Q_{O}(\epsilon)=\epsilon$ and
-$Q_{O}(w\cdot a)=\begin{cases}Q_{O}(w)\cdot m&\text{if }{q}^{\mathsf{T}}(w,a)=(deq,m)\text{ for some }m\in\mathit{Msgs}\\ Q_{O}(w)&\text{otherwise}\end{cases}$
+$Q_{O}(w\cdot a)=\begin{cases}Q_{O}(w)\cdot m&\text{if }w_{3}(w\cdot a)=w_{3}(w)\cdot(deq,m)\text{ for some }m\in\mathit{Msgs}\\ Q_{O}(w)&\text{otherwise}\end{cases}$
 This is the queue of messages that has passed through the device to the output buffer.
@@
 Assuming $Tspec(w)$
-$\text{If }WaitEnq(w)=0\text{ then }Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))$
-$\text{If }WaitEnq(w)=1\text{ then }Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))\cdot Inval(in(w))$
+$\text{If }WaitEnq(w)=0\text{ then }Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(w_{3}(w))$
+$\text{If }WaitEnq(w)=1\text{ then }Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(w_{3}(w))\cdot Inval(in(w))$
+In either case, $Q_{O}(w)$ is a prefix of $Q_{I}(w)$.
 The difference is that when $WaitEnq(w)=1$ the value in the input buffer is the last item that has shown up on the input queue, otherwise the input buffer can be ignored.
-It’s an immediate consequence of this theorem that $Q_{O}(w)$ is a prefix of $Q_{I}(w)$.
+This theorem also shows that the protocol works as long at $Tspec(w)$ remains true.
 The proof of 1 is by induction on the event sequence.
+The proof here is in too much detail, just to show it can be done completely.
 Initially, when $w=\epsilon$, all the queues are empty and $WaitEnq(\epsilon)=0$ so there is nothing to prove.
@@
 We assume $Tspec(w\cdot a)$ otherwise there is nothing to prove.
-There are four cases.
+There are four cases depending on the four possible values of $(WaitEnq(w),WaitEng(w\cdot a))$.
   1. 1.
@@
     By the definition of $WaitEnq$, one of the two toggles has to change thanks to $a$.
-If ${q}^{\mathsf{T}}(w,a)=(deq,m)$ that would change the $Last$ toggle but because $WaitEnq(w)=0$ ${q}^{\mathsf{T}}(w,a)\neq(deq,m)$ for any $m$ so it must be that $a=(input,m)$ for some $m$ and $Inb$ changes.
-By the inductive assumption and because $WaitEnq(w)=0$ $Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))$.
+Because $WaitEnq(w)=0)$, $w_{3}(w\cdot a)$ cannot add a $(deq,m)$ event that would change the $Last$ toggle.
+So the alternative must be that $a=(input,m)$ for some $m$.
+$m$ and $Inb$ changes.
+By the inductive assumption and because $WaitEnq(w)=0$ $Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(w_{3})$.
 Because $a=(input,m)$, $Q_{I}(w\cdot a)=Q_{I}(w)\cdot m$.
-By the definition of ${inv}^{\mathsf{T}}$ and because $a=(input,m)$, $Inval(inv(w\cdot a))=m$.
-We know $a\neq deq$ so ${q}^{\mathsf{T}}(w,a)\neq deq$ – so $Q(q(w))=Q(q(w\cdot a))$.
-This proves the first case.
+By the definition of $w_{1}$, it must be that $Inval(w\cdot a)=m$ We know $a\neq deq$ so $w_{3}(w\cdot a)$ cannot add deqdeq – so $Q(w_{3}(w))=Q(w_{3}(w\cdot a))$.
+It follows that $Q_{I}(w\cdot a)=Q_{O}(w\cdot a)\mathop{\mathtt{concat}}Q(w_{3}(w\cdot a)\cdot Inval(w_{1}(w\cdot a)$.
   2. 2.
@@
     In this case $a\neq(input,m)$ because otherwise $WaitEnq(w\cdot a)=1$ unless the $Last$ toggle flips too.
-But $Last$ cannot change because that needs ${q}^{\mathsf{T}}(w,a)=(enq,m)$ and that cannot be true if $WaitEnq(w)=0$.
-So $Q_{I}(w\cdot a)=Q_{I}(w))$.
-If $a=deq$ then if $Q(q(w))=\epsilon$ nothing happens to $Q_{O}$ or $Q$.
-If $Q(q(w))\neq\epsilon$ and $a=deq$, then ${q}^{\mathsf{T}}(w,a)=deq$ so $Q(q(w\cdot a)=Tail(Q(w))$ and $Q_{O}(w\cdot a)=Q_{O}(w)\cdot Head(Q(q(w)))$ but
-    $Q_{O}(w)\cdot Head(Q(q(w))\mathop{\mathtt{concat}}Tail(Q(q(w))$
+But $Last$ cannot change because $WaitEnq(w)=0$ means that $w_{3}(w\cdot a)$ does not add an $(enq,m)$ which is required to toggle $Last$.
+Because $a\neq(input,m)$ it follows that $Q_{I}(w\cdot a))=Q_{I}(w))$.
+If $a=deq$ then if $Q(w_{3}(w))=\epsilon$ nothing happens to $Q_{I}$, $Q_{O}$ or $Q$.
+If $Q(w_{3}(w))\neq\epsilon$ and $a=deq$, then $w_{3}(w\cdot a)=w_{3}(w)\cdot deq$ so $Q(w_{3}(w\cdot a)=Tail(Q(w_{3}(w)))$ and $Q_{O}(w\cdot a)=Q_{O}(w)\cdot Head(Q(w_{3}(w)))$ but
+    $Q_{O}(w)\cdot Head(Q(w_{3}(w))\mathop{\mathtt{concat}}Tail(Q(w_{3}(w))$
     must be the same as
-    $Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w)).$
-    Otherwise ${q}^{\mathsf{T}}(w,a)=\epsilon$ That takes care of this case.
+    $Q_{O}(w)\mathop{\mathtt{concat}}Q(w_{3}(w)).$
+    Otherwise $w_{3}(w\cdot a)$ adds $\epsilon$ and nothing changes. That takes care of this case.
   3. 3.
@@
     $WaitEnq(w)=1\text{ and }WaitEnq(w\cdot a)=1$
-    Because $WaitEnq$ remains true, ${q}^{\mathsf{T}}(w,a)\neq(enq,m)$ for any $m$ so no new elements are added to $Q(q(w\cdot a))$.
-Because we assume $Tspec(w\cdot a)$, $WaitEnq(w)*SinceI(w)<t_{0}$ so $a$ is not an input event so ${inv}^{\mathsf{T}}(w,a)=\epsilon$ and $Inval(inv(w\cdot a))=Inval(inv(w))$.
-If $a\neq deq$, it must be that ${q}^{\mathsf{T}}(w,a)=\epsilon$ so $Q(q(w\cdot a))=Q(q(w))$ when $a\neq deq$ and similarly $Q_{O}(w\cdot a)=Q_{O}(w)$ because ${ov}^{\mathsf{T}}(w,a)=\epsilon$.
-If $a=deq$ then than $Q(q(w\cdot a))=Tail(Q(q(w)))$ and $Q_{O}(w\cdot a)=Q_{O}(w)\cdot Head(Q(q(w)))$ so the constructed queue is unchanged.
+    Because $WaitEnq$ remains true, $w_{3}(w\cdot a)$ does not add $(enq,m)$ for any $m$ so no new elements are added to $Q(w_{3}(w\cdot a))$.
+Because we assume $Tspec(w\cdot a)$, $WaitEnq(w)*SinceI(w)<t_{0}$ so $a$ is not an $(input,m)$ for any $m$ and thus $w_{1}(w\cdot a)=w_{1}(w)$ and $Inval(inv(w\cdot a))=Inval(inv(w))$.
+If $a\neq deq$, it must be that $w_{3}(w\cdot a)$ adds $\epsilon$ so $Q(w_{3}(w\cdot a))=Q(w_{3}(w))$ and similarly $Q_{O}(w\cdot a)=Q_{O}(w)$ because $w_{4}(w\cdot a)i=w_{4}(w)$.
+If $a=deq$ then then $Q(w_{3}(w\cdot a))=Tail(Q(w_{3}(w)))$ and $Q_{O}(w\cdot a)=Q_{O}(w)\cdot Head(Q(w_{3}(w)))$ so the constructed queue is unchanged.
     For any other $a$ nothing changes in the outputs of any components and we are done with this case.
@@
     Since $WaitEnq(w)=1$
-    $Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(q(w))\cdot Inv(inv(w).$
+    $Q_{I}(w)=Q_{O}(w)\mathop{\mathtt{concat}}Q(w_{3}(w))\cdot Inv(inv(w).$
     Either $SinceI(w)<t_{0}$ or $TSpec(w)=0$.
 We’re assuming the second is false, so by the same reasoning $a\neq(input,m)$ (or else $Tspec(w\cdot a)=0$).
-That means that $q(w,a)=(enq,Inv(inv(w)))$ to make $WaitEnq(w\cdot a)=0$.
+That means that $w_{3}(w\cdot a)$ adds $(enq,Inv(w_{1}(w)))$ to make $WaitEnq(w\cdot a)=0$.
 As a result
-    $Q(q(w\cdot a))=Q(q(w))\cdot Inv(inv(w)).$
+    $Q(w_{3}(w\cdot a))=Q(w_{3}(w))\cdot Inv(w_{1}(w)).$
     And the rest follows.
@@
 5 Discussion
-When Professor Krithivasan Ramamritham gave me temporal logic papers to read[12] at the start of my graduate studies in the 1980s, I had just finished working on a commerical distributed operating system[2] and was frustrated by how often the developers had found design errors only during test (after writing enormous quantities of code).
-Temporal logic seemed like a solution, but I was not unique in finding that it was not a completely satisfactory answer.
-Treatment of composition and concurrency, the inflexibility of the temporal quantifiers, the ad hoc nature of the formal semantics were among the issues that made it difficult to get a handle on the complexity of actual operating system design.
-Much of the methods used here and in [14] were present in a 1991 paper [13], which included Moore machines, automata products for composition and concurrency and even a modal primitive recursive specification language.
+When Krithivasan Ramamritham gave me temporal logic papers to read[14] at the start of my graduate studies in the 1980s, I had just finished working on a commercial distributed operating system[2] and was frustrated by how often developers found design errors only during test after writing enormous quantities of code.
+Temporal logic seemed like a solution, but I was not unique in concluding is was not a satisfactory answer after attempting to apply it to real systems.
+Lamport and Abadi argued (in 1994) that ”It has long been known that the methods originally developed for shared memory multiprocessing apply equally well to distributed systems” (p.
+1543) but it’s not clear that they do apply well at all.
+Much of the complexity of the Abadi and Lamport paper comes from the assumption that there is a “fair scheduler” somewhere deciding which nondeterministic action to activate – for example, to advance the time variable.
+This view is reflected in the use of infinite sequences of assignment maps (maps assigning values to formal variables ) as formal semantics.
+Infinite sequences of assignment maps that must be composed by interleaving and where semantics depends a scheduler that exists outside the quantifiers inflexibility of the temporal quantifiers, the ad hoc nature of the formal semantics were among the issues that made it difficult to get a handle
+Some temporal logic variants use infinite sequences of assignment maps (maps from propositions or variables to values) as formal semantics.
+modal and temporal logics can define state machines or use state machines as semantic models.
+Kripke’s initial modal logic paper[8] makes this clear although he did not use the terminolgy of state machines.
+Finite state machines, of a kind, can be constructed from logical assertions to serve as models for temporal logic as well[4, 3]:
+Temporal logic formulae are interpreted over a given finite state graph, also called a (Kripke) structure, $M$ comprised of a set $S$ of states, a total binary transition relation $R\subseteq S\times S$, and a labelling $L$ of states with atomic facts (propositions) true there.
+– Edmond Clarke [5]
+These are non-deterministic machines of a sort with unlabeled state transitions which may be controlled by logical predicates.
+States are assignment maps assigning values to formal symbols, and composition is by interleaving and essentially taking the union of the assigment maps assuming the domains are disjoint.
+The view is still derived from concurrent processes in a timeshared single processor computer system and there is clear model of composition.
+Methods used here and in [17] were present in a 1991 paper [16], which included Moore machines, automata products for composition and concurrency and even a modal primitive recursive specification language.
 But that work was in the tradition of computational formal logic and this is not.
@@
 Whatever the merits of that approach, discarding the apparatus of formal logic has allowed solution of a number of problems that blocked progress in the earlier method and greatly simplified specifications.
-The quote from Von Neumann that starts this paper is from a, not surprisingly, enormously perceptive paper that among other things previews algorithmic complexity theory.
+The quote from Von Neumann that starts this paper is from an enormously perceptive paper (not unusual for the author) that among other things previews algorithmic complexity theory.
 The full paragraph in which that quote is found has an interesting critique which I have tried to learn from as I returned to this work after a long absence.
@@
 $Tail(\epsilon)=\epsilon,Tail(rappend(w,a))=w$
-Appendix B Appendix: Sequential functions
-This is a short summary of results from [14].
-The standard presentation of a Moore machine is $(A,X,S,s_{0},\delta,\lambda)$ where $s_{0}\in S$ is the start state, $\delta:S\times A\to S$ and $\lambda:S\to X$.
-Primitive recursive sequential functions are reformulations to make the state set implicit and the recursion explicit.
-$D(\epsilon)=c,\ D(w\cdot a)=g(D(w),a)\text{ and }M(w)=L(D(w))$
-Such a map can be transformed into a conventional (possibly infinite state) Moore machine by letting the state set be the set $\{D(w):w\in A^{*}\}$.
-A sequential function is any map $f:A^{*}\to X$ where $A$ is a set of events, $A^{*}$ are the finite sequences over $A$ and $X$ is the set of outputs.
-Sequential functions define possibly infinite state Moore machines via the usual congruence on sequences.
-The primitive recursive sequence functions are a proper subset with some useful properties including $f(w)$ must be defined if $f(w\cdot a)$ is defined.
-The finite state sequential functions are a proper subset of the primitive recursive sequential functions that correspond to finite state Moore machines.
-Specifications of real devices and programs tend to use primitive recursive and finite state primitive recursive sequential functions without any special attention because properties of real systems are constructive and real systems are finite state machines anyway.
-Primitive recursive sequential functions are constructed by a finite number of applications of the following rules.
-  1. 1.
-    Constant. If $c$ is a constant $f(w)=c$ defines a sequential function.
-  2. 2.
-    Constant map. If $g$ is a non-state dependent map, $f(w,x)=g(x)$ defines a sequential function.
-  3. 3.
-    Primitive Recursion on Sequences. If $g$ is a non-state dependent map, $c$ is a constant, $\epsilon$ is the empty sequence
-    $f(\epsilon)=c,\ f(w\cdot a)=g(a,f(w)))$
-    defines a sequential function. If $g$ has a finite image $f$ is finite state.
-  4. 4.
-    Simple Composition. If $f_{1}$ is a sequential function, and $h(x)$ is an ordinary map
-    $f(w)=h(f_{1}(w))$
-    defines a sequential function.
-As with the arithmetic primitive recursive functions[9], many additional forms can be defined without changing the class of primitive recursive sequential functions.
-In particular multiple arguments $f(w,x_{1},\dots x_{n})$ are fine, even with some arguments that are also state dependent $f(w)=(f(v_{1}(w),\dots f_{n}(w)))$.
-An analog of arithmetic simultaneous recursion allows definition of composite system which correspond to automata products and specify concurrent/parallel systems of machines.
-  • General Product
-    Given sequential functions $f_{1},\dots f_{n}$, each $f_{i}:A_{i}^{*}\to X_{i}$ and appropriate maps $g_{i}:A\times X\to A_{i}^{*}$ (not sequence dependent) where $X=X_{1}\times\dots X_{n}$,
-    $f(w)=(f_{1}(u_{1}(w)),\dots f_{n}(u_{n}(w)))$
-    $u_{i}(\epsilon)=\epsilon,\text{ and }u_{i}(w\cdot a)=u_{i}(w)\mathop{\mathtt{concat}}g(a,i,v(w))$
-As shown in [14]: If each of the $f_{i}$ is primitive recursive, then $f$ is primitive recursive. If each $f_{i}$ is finite state and $A$ is finite then $f$ is finite state.